Rewrite the given equation in slope-intercept form:
- 3x + y = 5
Please show how you arrived at answer. Thank you.

Answer by pragya.ban@gmail.com(32)

- 3x + y = 5

Slope intercept form is nothing but y=mx+b (where m is the slope and b is the y intercept)
The first step to solving this problem is to isolate y on one side and everything else on another.
To do so, you have to add x on both sides of of the equation. It will then look like this:

-3x+3x+y=3x+5

-3x and +3x cancel each other out/equal to zero so you get:

y=3x+5

As you can see, this matches the y=mx+b. (m=3 and b=5)

Good luck!